1.Order the following elements from most moles in 1 kg at the top to fewest moles in 1 kg at the bottom.

Lead, radium,thorium,uranium.

2.For what reasons would a nucleus undergo radioactive decay? Give at least two reasons.

3.In what way are nuclear fusion and fission similar? How are these two processes different?
Answer #1

1. To find the number of moles in 1 kg of each element, we need to know their molar masses.

Molar mass of Pb = 207 g/mol

Molar mass of Ra = 226 g/mol

Molar mass of Th = 232 g/mol

Molar mass of U = 238 g/mol

Number of moles = Given weight / Molecular weight ---------------equation (i)

The weight given is common for all i.e 1 kg or 1000 g.

Hence substituting the values for molar mass and given mass of each element in equation (i) we have,

Number of moles of Pb = 1000/ 207 = 4.83

Number of moles of Ra = 1000/ 226 = 4.42

Number of moles of Th = 1000/ 207 = 4.31

Number of moles of U = 1000/ 238 = 4.20

Hence the order of the given elements from maximum number of moles to minimum is:

Pb>Ra>Th>U

2. The stability of a nucleus depends upon the number of neutrons and protons present in it. When the number of number of neutrons to the number of protons ratio (n/p) is equal to 1, a nucleus is stable. But if n/p > 1 or if n/p < 1, the nucleus becomes unstable. This is because with the increase in the number of protons, the force of repulsion between them would increase which would lower the stability of the nucleus. Such an unstable nuclei would attain stability by undergoing radioactive decay to produce a stable nucleus.

a. If the n/p > 1, i.e if the number of neutrons is more than the number of protons, the nucleus disintegrates in such a way that one of their neutrons get converted to a proton such that the neutron and proton number becomes equal.

b. If the n/p < 1, i.e if the number of protons is more than the number of neutrons, the nucleus disintegrates in such a way that one of their protons get converted to a neutron such that the neutron and proton number becomes equal.

3. Similarities: Both nuclear fusion and fission reactions produce a large amount of energy.

Both nuclear fusion and fission reactions are nuclear reactions i.e they involve a change in the nucleus of the atom.

Differences: Fusion is a process in which two lighter nuclides fuse to form a heavier nuclide. Fission is a process in which a heavier nuclide on being bombarded by some high speed particle (example neutrons) splits up into two lighter nuclide.

Nuclear fission is easier to control compared to Nuclear fusion.

Fission : 238 U + 1n0\rightarrow141 Ba + 92 Kr + 3 1n0

Fusion : 2 H + 2H \rightarrow4He